# Energy Conditions and Quantum Information

### v1

@@ 2.6.1 Black hole mechanics and energy conditions
{itembox}
[l]Proposition 2.6.1 If the NEC holds, then a closed trapped surface $\cal T$}, if it exists, cannot intersect $J^{-}(\mathscr{I}^{+})$.
  • Propositoin 9.2.1 HE73 .
@@
The proof goes roughly as follows. To derive a contradiction, suppose ${\cal T}\in J^{-}(\mathscr{I}^{+})$. Then, a null geodesic generator $\mu$ of $\partial J^{+}({\cal T})$ has an intersection with $\mathscr{I}^{+}$ in $\tilde{M}$. This implies on one hand that $\mu$ is future complete, and on the other hand that $\mu$ has the past endpoint at $\cal T$}. Then, a null geodesic congruence of $\mu$ orthogonal to $\cal T$} must have a negative expansion $\theta_{+}<0$ at $\cal T$}, since $\cal T$} is a closed trapped surface. Then, by Theorem 2.4.2, $\mu$ must have a conjugate point to $\cal T$} in the future of $\cal T$}. This, however, contradicts the fact that $\mu$ is the generator of the achronal hypersurface $\partial J^{+}({\cal T})$ (Proposition 2.5.1). $\Box$
@@
{itembox}
[l]Theorem 2.6.2 (Area theorem: The second law) Under the NEC, the cross-sectional area of the event horizon never decreases toward future.
  • Propositoin 9.2.1 HE73 , Theorem 12.2.6 Wald84 .
@@
An asymptotically flat spacetime is said to be stationary, if there is a Killing vector field, $t^{\mu}$ whose orbits are complete (i.e., $t^{\mu}$ generates an isometry group) and timelike, at least near infinity $\mathscr{I}$. Similarly, a spacetime is said to be axisymmetric, if there exists a Killing vector field $\varphi^{\mu}$ whose orbits are closed, spacelike curves. We also assume that the two Killing vector fields $t^{\mu}$ and $\varphi^{\mu}$ commute. An important consequence is that: {itembox}[l]Theorem 2.6.3 (Rigidity theorem) If an asymptotically flat, analytic, stationary spacetime solving the electrovacuum Einstein equations contains a black hole, then the event horizon of the black hole must be a Killing horizon. Furthermore, if a stationary black hole is rotating, it must also be axisymmetric.
@@
Although the surface gravity $\kappa$, defined as above, is a local notion, the following result holds: {itembox}[l]Theorem 2.6.4 (The zero-th law) The surface gravity $\kappa$ of a stationary black hole must be a constant over its event horizon.
@@
$\delta M=\oint_{\sigma}{\rm d}S\int_{0}^{\infty}\!\!\!\!\!\!{\rm d}\lambda\;T_{\mu\nu}k^{\mu}t^{\nu}\,,\quad\delta J=-\oint_{\sigma}{\rm d}S\int_{0}^{\infty}\!\!\!\!\!\!{\rm d}\lambda\;T_{\mu\nu}k^{\mu}\varphi^{\nu}\,$, (39)
where $\sigma$ denotes a horizon cross-section with the area element ${\rm d}S$, and $k^{\mu}={\rm d}x^{\mu}/{\rm d}\lambda$ the tangent to the horizon null geodesic generator with $\lambda$ being an affine parameter so that $k^{\mu}=\chi^{\mu}{\rm d}v/{\rm d}\lambda=\chi^{\mu}/(\kappa\lambda)=(t^{\mu}+\Omega_{H}\varphi^{\mu})/(\kappa\lambda)$. At this stage, we regard $T_{\mu\nu}$ as a small perturbation, which generates non-vanishing expansion and shear via the Raychaudhur equation. In the linear order of perturbation, neglecting the quadratic terms of the expansion and shear, we can write the Raychaudhuri equation for the horizon null geodesic generators as
$\dfrac{{\rm d}\theta}{{\rm d}\lambda}\simeq-8\pi GT_{\mu\nu}k^{\mu}k^{\nu}=-\dfrac{8\pi G}{\kappa\lambda}T_{\mu\nu}(t^{\mu}+\Omega_{H}\varphi^{\mu})k^{\nu}\,.$ (40)
Then, by integrating the Raychaudhuri equation along the horizon generators [^9] By multiplying $\kappa\lambda$ on both side of (40), and taking integration over $\sigma$ and along the null generator, the right-hand side immediately yields $8\pi G(\delta M-\Omega_{H}\delta J)$, while the left-hand side gives $\oint_{\sigma}{\rm d}S\int_{0}^{\infty}\!\!\!\!\!\!{\rm d}\lambda\,(\kappa\lambda)\,\dfrac{{\rm d}\theta}{{\rm d}\lambda}=\kappa\oint_{\sigma}{\rm d}S\int_{0}^{\infty}\!\!\!\!\!\!{\rm d}\lambda\left[\dfrac{{\rm d}(\lambda\theta)}{{\rm d}\lambda}-\theta\right]=-\kappa\oint_{\sigma}{\rm d}S\int_{0}^{\infty}\!\!\!\!\!\!{\rm d}\lambda\,\theta=-\kappa\delta A\,$, (41) where we have used the fact that before and sufficiently after one has dropped matter fields into the horizon, $\theta=0$ as a stationary state. , as well as horizon cross-section $\sigma$, we have {itembox}[l]Theorem 2.6.5 (The first law)
$\dfrac{\kappa}{8\pi G}\delta A=\delta M-\Omega_{H}\delta J$ (42)
@@ 2.6.2 Topology and energy conditions
To discuss horizon topology, we first recall the notion of stably marginally outer trapped surface (MOTS), which we have introduced in the previous section, as any $2$-sphere on the future event horizon of the Schwarzschild black hole. To define the notion of MOTS in more general setting, consider a closed, connected $2$-surface ${\cal T}_{0}$. Let $k^{\mu}$ be the tangent vector field for an out-going101010 In general, the notion of “out-going” or “outward” direction is not uniquely given. For an asymptotically flat spacetime with a single asymptotic region, if ${\cal T}_{0}$ separates a Cauchy surface $\Sigma$ into two disconnected parts $\Sigma_{\rm in}$ and $\Sigma_{\rm out}$, then “out-going” direction is defined as the one, say $\Sigma_{\rm out}$, that contains the asymptotic region. null geodesic congruence orthogonally emanating from ${\cal T}_{0}$, and $\theta_{+}$ be the associated expansion. Let $\ell^{\mu}$ be a past-directed, outward null vector field on ${\cal T}_{0}$. If $\theta_{+}=0$ and $\ell^{\nu}\nabla_{\nu}\theta_{+}\geqslant 0$ on ${\cal T}_{0}$, then ${\cal T}_{0}$ is said to be a MOTS. The following theorem determines the topology of MOTS.
{itembox}
[l]Theorem 2.6.6 (Topology of horizon cross-section) Under the DEC, the horizon cross-section ${\cal T}_{0}$ of a stationary black hole must be topologically $2$-sphere.
  • Propositoin 9.3.2 HE73 .
@@
$\oint_{{\cal T}_{0}}\!\!{\rm d}S\,{\cal R}\geqslant 8\pi T_{\mu\nu}k^{\mu}(-\ell^{\nu})\geqslant 0\,$, (47)
@@
{itembox}
[l]Theorem 2.6.7 (Topology and singularity) Let $(M,g_{\mu\nu})$ be a globally hyperbolic, asymptotically flat spacetime, on which the NEC holds. Suppose that $M$ contains a non-simply connected Cauchy surface $\Sigma$. Then, $M$ must be geodesically incomplete.
@@
Can a distant observer actually detect this wormhole (handle) structure within that bounded region inside $\sigma$? For example, we would attempt to send a probe rocket into this bounded region so that the rocket would orbit the handle, become snagged by its non-trivial topological structure, and then return to a distant location to report its findings. Again under the NEC, we have the following theorem. {itembox}[l]Theorem 2.6.8 (Topological censorship) Let $(M,g_{\mu\nu})$ be a globally hyperbolic, asymptotically flat spacetime. If the NEC holds in $M$, then any causal curve $\gamma$ from a point on $\mathscr{I}^{-}$ to a point on $\mathscr{I}^{+}$ can be continuously deformed to a timelike curve $\gamma_{0}$ which is contained in a simply connected neighborhood of $\mathscr{I}^{+}\cup\mathscr{I}^{-}$.
@@
Applying Theorem 2.6.8 to $\rm DOC$}, one can find the following constraints on the horizon topology:
{itembox}
[l]Theorem 2.6.9 Suppose $(M,g_{\mu\nu})$ is a stationary, asymptotically flat spacetime containing a single asymptotically flat region, whose $\rm DOC$} is globally hyperbolic. Suppose the NEC holds. Then, (i) $\rm DOC$} is simply connected, and (ii) each connected component of the horizon cross-section is homeomorphic to a $2$-sphere.
@@
{itembox}
[l]Proposition 2.6.10 (No traversable wormhole) The traversable wormhole structure (TW) must violate the NEC.
@@ 2.7 Averaged energy conditions
So far we have reviewed the pointwise (locally defined) energy conditions (WEC, DEC, SEC, NEC) and their roles in classical general relativity. However, these conditions can be violated even within classical frameworks. For example, consider a scalar field $\phi$ with potential $V(\phi)$ with the energy-momentum tensor (7) in a spatially homogeneous and isotropic (FLRW) universe. Let $t^{\mu}=(\partial/\partial t)^{\mu}$ be a unit timelike vector orthogonal to any homogeneous and isotropic hypersurface. Then, due to the symmetry, $\phi$ can only be dependent on $t$, and we have $\left(T^{(\phi)}_{\mu\nu}-({1}/{2})T^{(\phi)}{}^{\lambda}{}_{\lambda}g_{\mu\nu}\right)t^{\mu}t^{\nu}=\dot{\phi}^{2}-V(\phi)$. Thus, if $\dot{\phi}^{2}<V$, which can occur for e.g., a massive free scalar field, or inflation/dark energy models, the SEC fails to hold.
@@
In the semiclassical approach, quantum field effects can be incorporated in the expectation value of the renormalized energy-momentum tensor $\langle T_{\mu\nu}\rangle$, which replaces the classical counterpart in the right-hand side of the Einstein equations (1). There are a number of known examples in which the expectation value $\langle T_{\mu\nu}\rangle$ violates the locally defined energy conditions, even when their classical counterparts do not violate the locally defined energy conditions. In fact, Casimir effects violate all local energy conditions Roman:1986tp . For example, consider two parallel reflecting planes orthogonal to $z$-axis in Minkowski spacetime. Then the vacuum expectation value of the energy-momentum tensor for a scalar field between the two planes takes the form of $\langle 0|T_{\mu\nu}|0\rangle=-(C^{2}/L^{4})\mathrm{diag}(1,-1,-1,3)$ with $C^{2}$ being a positive constant and $L$ the separation distance between the two planes Birrell:1982ix , for which all the four locally defined energy conditions are violated.
@@
While the pointwise energy conditions may fail in some restricted local regions, they may still hold in the rest of the universe. This observation suggests that non-local forms of energy conditions, which are defined by taking a suitable average over spacetime, may be more robust under many circumstances of interest. For example, in the case of Casimir effect above, although between the two planes, the NEC is violated along $z$-axis, since the two planes supply positive energy, if we integrate $\langle 0|T_{\mu\nu}k^{\mu}k^{\nu}|0\rangle$ along $z$-axis long enough passing through the two planes, we can have a positive energy value. In what follows, we consider generalizations of the energy conditions to the averaged forms. As we have seen in Sec. 2.4, 2.5, and 2.6, the energy conditions play a key role mainly through the focusing theorems, we will focus on the averaged SEC and NEC. As a weaker condition than SEC, the averaged strong-energy condition (ASEC) is simply formulated as follows.
{itembox}
@@
[l]Definition 2.7.1 (ASEC) For every inextendible timelike geodesic curve with proper time $\tau$ and the tangent $\xi^{\mu}$, the integral along the timelike geodesic,
$\int{\rm d}\tau\left(T_{\mu\nu}-\dfrac{1}{2}Tg_{\mu\nu}\right)\xi^{\mu}\xi^{\nu}\geqslant 0\,.$ (50)
@@
{itembox}
[l]Definition 2.7.2 (ANEC) For every inextendible null geodesic curve with affine parameter $\lambda$ and the tangent $k^{\mu}$, the integral along the null geodesic,
$\int{\rm d}\lambda\,T_{\mu\nu}k^{\mu}k^{\nu}\geqslant 0\,.$ (52)

### v2

@@ 2.6.1 Black hole mechanics and energy conditions
  Proposition 2.6.1   If the NEC holds, then a closed trapped surface $\cal T$}, if it exists, cannot intersect $J^{-}(\mathscr{I}^{+})$.
  • Proposition 9.2.1 HE73 .
@@
The proof goes roughly as follows. To derive a contradiction, suppose ${\cal T}\cap J^{-}(\mathscr{I}^{+})\neq\emptyset$. Then, a null geodesic generator $\mu$ of $\partial J^{+}({\cal T})$ has an intersection with $\mathscr{I}^{+}$ in $\tilde{M}$. This implies on one hand that $\mu$ is future complete, and on the other hand that $\mu$ has the past endpoint at $\cal T$}. Then, a null geodesic congruence of $\mu$ orthogonal to $\cal T$} must have a negative expansion $\theta_{+}<0$ at $\cal T$}, since $\cal T$} is a closed trapped surface. Then, by Theorem 2.4.2, $\mu$ must have a conjugate point to $\cal T$} in the future of $\cal T$}. This, however, contradicts the fact that $\mu$ is the generator of the achronal hypersurface $\partial J^{+}({\cal T})$ (Proposition 2.5.1). $\Box$
@@
  Theorem 2.6.2 (Area theorem: The second law)   Under the NEC, the cross-sectional area of the event horizon never decreases toward future.
  • Proposition 9.2.1 HE73 , Theorem 12.2.6 Wald84 .
@@
An asymptotically flat spacetime is said to be stationary, if there is a Killing vector field, $t^{\mu}$ whose orbits are complete (i.e., $t^{\mu}$ generates an isometry group) and timelike, at least near infinity $\mathscr{I}$. Similarly, a spacetime is said to be axisymmetric, if there exists a Killing vector field $\varphi^{\mu}$ whose orbits are closed, spacelike curves. We also assume that the two Killing vector fields $t^{\mu}$ and $\varphi^{\mu}$ commute. An important consequence is that:
  Theorem 2.6.3 (Rigidity theorem)   If an asymptotically flat, analytic, stationary spacetime solving the electrovacuum Einstein equations contains a black hole, then the event horizon of the black hole must be a Killing horizon. Furthermore, if a stationary black hole is rotating, it must also be axisymmetric.
@@
Although the surface gravity $\kappa$, defined as above, is a local notion, the following result holds:
  Theorem 2.6.4 (The zero-th law)   The surface gravity $\kappa$ of a stationary black hole must be a constant over its event horizon.
@@
$\delta M$ $=$ $\oint_{\sigma}{\rm d}S\int_{0}^{\infty}\!\!\!\!\!\!{\rm d}\lambda\;T_{\mu\nu}k^{\mu}t^{\nu}\,$,
$\delta J$ $=$ $-\oint_{\sigma}{\rm d}S\int_{0}^{\infty}\!\!\!\!\!\!{\rm d}\lambda\;T_{\mu\nu}k^{\mu}\varphi^{\nu}\,$, (39)
where $\sigma$ denotes a horizon cross-section with the area element ${\rm d}S$, and $k^{\mu}={\rm d}x^{\mu}/{\rm d}\lambda$ the tangent to the horizon null geodesic generator with $\lambda$ being an affine parameter so that $k^{\mu}=\chi^{\mu}{\rm d}v/{\rm d}\lambda=\chi^{\mu}/(\kappa\lambda)=(t^{\mu}+\Omega_{H}\varphi^{\mu})/(\kappa\lambda)$. At this stage, we regard $T_{\mu\nu}$ as a small perturbation, which generates non-vanishing expansion and shear via the Raychaudhuri equation. In the linear order of perturbation, neglecting the quadratic terms of the expansion and shear, we can write the Raychaudhuri equation for the horizon null geodesic generators as
$\dfrac{{\rm d}\theta}{{\rm d}\lambda}$ $\simeq$ $-8\pi GT_{\mu\nu}k^{\mu}k^{\nu}$ (40)
$=$ $-\dfrac{8\pi G}{\kappa\lambda}T_{\mu\nu}(t^{\mu}+\Omega_{H}\varphi^{\mu})k^{\nu}\,.$
Then, by integrating the Raychaudhuri equation along the horizon generators [^10] By multiplying $\kappa\lambda$ on both side of (40), and taking integration over $\sigma$ and along the null generator, the right-hand side immediately yields $8\pi G(\delta M-\Omega_{H}\delta J)$, while the left-hand side gives $\oint_{\sigma}{\rm d}S\int_{0}^{\infty}\!\!\!\!\!\!{\rm d}\lambda\,(\kappa\lambda)\,\dfrac{{\rm d}\theta}{{\rm d}\lambda}$ $=$ $\kappa\oint_{\sigma}{\rm d}S\int_{0}^{\infty}\!\!\!\!\!\!{\rm d}\lambda\left[\dfrac{{\rm d}(\lambda\theta)}{{\rm d}\lambda}-\theta\right]$ (41) $=$ $-\kappa\oint_{\sigma}{\rm d}S\int_{0}^{\infty}\!\!\!\!\!\!{\rm d}\lambda\,\theta$ $=$ $-\kappa\delta A\,$, where we have used the fact that before and sufficiently after one has dropped matter fields into the horizon, $\theta=0$ as a stationary state. , as well as horizon cross-section $\sigma$, we have
  Theorem 2.6.5 (The first law)   $\dfrac{\kappa}{8\pi G}\delta A=\delta M-\Omega_{H}\delta J$ (42)
@@ 2.6.2 Topology and energy conditions
To discuss horizon topology, we first recall the notion of stably marginally outer trapped surface (MOTS), which we have introduced in the previous section, as any $2$-sphere on the future event horizon of the Schwarzschild black hole. To define the notion of MOTS in more general setting, consider a closed, connected $2$-surface ${\cal T}_{0}$. Let $k^{\mu}$ be the tangent vector field for an out-going [^11] In general, the notion of “out-going” or “outward” direction is not uniquely given. For an asymptotically flat spacetime with a single asymptotic region, if ${\cal T}_{0}$ separates a Cauchy surface $\Sigma$ into two disconnected parts $\Sigma_{\rm in}$ and $\Sigma_{\rm out}$, then “out-going” direction is defined as the one, say $\Sigma_{\rm out}$, that contains the asymptotic region. null geodesic congruence orthogonally emanating from ${\cal T}_{0}$, and $\theta_{+}$ be the associated expansion. Let $\ell^{\mu}$ be a past-directed, outward null vector field on ${\cal T}_{0}$. If $\theta_{+}=0$ and $\ell^{\nu}\nabla_{\nu}\theta_{+}\geqslant 0$ on ${\cal T}_{0}$, then ${\cal T}_{0}$ is said to be a MOTS. The following theorem determines the topology of MOTS.
  Theorem 2.6.6 (Topology of horizon cross-section)   Under the DEC, the horizon cross-section ${\cal T}_{0}$ of a stationary black hole must be topologically $2$-sphere.
  • Proposition 9.3.2 HE73 .
@@
$\oint_{{\cal T}_{0}}\!\!{\rm d}S\,{\cal R}\geqslant 16\pi G\oint_{{\cal T}_{0}}\!\!{\rm d}S\,T_{\mu\nu}k^{\mu}(-\ell^{\nu})\geqslant 0\,$, (47)
@@
  Theorem 2.6.7 (Topology and singularity)   Let $(M,g_{\mu\nu})$ be a globally hyperbolic, asymptotically flat spacetime, on which the NEC holds. Suppose that $M$ contains a non-simply connected Cauchy surface $\Sigma$. Then, $M$ must be geodesically incomplete.
@@
Can a distant observer actually detect this wormhole (handle) structure within that bounded region inside $\sigma$? For example, we would attempt to send a probe rocket into this bounded region so that the rocket would orbit the handle, become snagged by its non-trivial topological structure, and then return to a distant location to report its findings. Again under the NEC, we have the following theorem.
  Theorem 2.6.8 (Topological censorship)   Let $(M,g_{\mu\nu})$ be a globally hyperbolic, asymptotically flat spacetime. If the NEC holds in $M$, then any causal curve $\gamma$ from a point on $\mathscr{I}^{-}$ to a point on $\mathscr{I}^{+}$ can be continuously deformed to a timelike curve $\gamma_{0}$ which is contained in a simply connected neighborhood of $\mathscr{I}^{+}\cup\mathscr{I}^{-}$.
@@
Applying Theorem 2.6.8 to $\rm DOC$}, one can find the following constraints on the horizon topology.
  Theorem 2.6.9   Suppose $(M,g_{\mu\nu})$ is a stationary, asymptotically flat spacetime containing a single asymptotically flat region, whose $\rm DOC$} is globally hyperbolic. Suppose the NEC holds. Then, (i) $\rm DOC$} is simply connected, and (ii) each connected component of the horizon cross-section is homeomorphic to a $2$-sphere.
@@
  Proposition 2.6.10 (No traversable wormhole)   The traversable wormhole structure (TW) must violate the NEC.
@@ 2.7 Averaged energy conditions
So far we have reviewed the pointwise (locally defined) energy conditions (WEC, DEC, SEC, NEC) and their roles in classical general relativity. However, these conditions can be violated even within classical frameworks. For example, consider a scalar field $\phi$ with potential $V(\phi)$ with the energy-momentum tensor (7) in a spatially homogeneous and isotropic (FLRW) universe. Let $t^{\mu}=(\partial/\partial t)^{\mu}$ be a unit timelike vector orthogonal to any homogeneous and isotropic hypersurface. Furthermore, let us assume that $\phi$ depends only on $t$ in accord with the FLRW symmetry. Then we have $\left(T^{(\phi)}_{\mu\nu}-({1}/{2})T^{(\phi)}{}^{\lambda}{}_{\lambda}g_{\mu\nu}\right)t^{\mu}t^{\nu}=\dot{\phi}^{2}-V(\phi)$. Thus, if $\dot{\phi}^{2}<V$, which can occur for e.g., a massive free scalar field, or inflation/dark energy models, the SEC fails to hold.
@@
In the semiclassical approach, quantum field effects can be incorporated in the expectation value of the renormalized energy-momentum tensor $\langle T_{\mu\nu}\rangle$, which replaces the classical counterpart in the right-hand side of the Einstein equations (1). There are a number of known examples in which the expectation value $\langle T_{\mu\nu}\rangle$ violates the locally defined energy conditions, even when their classical counterparts do not violate the locally defined energy conditions. In fact, Casimir effects violate all locally defined (pointwise) energy conditions Roman:1986tp . For example, consider two parallel reflecting planes orthogonal to $z$-axis in Minkowski spacetime. Then the vacuum expectation value of the energy-momentum tensor for a scalar field between the two planes takes the form of $\langle 0|T_{\mu\nu}|0\rangle=-(C^{2}/L^{4})\mathrm{diag}(1,-1,-1,3)$ with $C^{2}$ being a positive constant and $L$ the separation distance between the two planes Birrell:1982ix , for which all the four locally defined energy conditions are violated. See also, e.g., Fewster Fewster:2012yh and Section 10.2 of Lobo:2017cay (and references therein) for a more elaborated explanation that all pointwise energy conditions can be violated in quantum field theory.
@@
While the pointwise energy conditions may fail in some restricted local regions, they may still hold in the rest of the universe. This observation suggests that non-local forms of energy conditions, which are defined by taking a suitable average over spacetime, may be more robust under many circumstances of interest. For example, in the case of Casimir effect above, although between the two planes, the NEC is violated along $z$-axis, since the two planes supply positive energy, if we integrate $\langle 0|T_{\mu\nu}k^{\mu}k^{\nu}|0\rangle$ along $z$-axis long enough passing through the two planes, we can have a positive energy value. In this respect, see also Ford:1978qya for a discussion concerning the second law of thermodynamics and averaged energy conditions. In what follows, we consider generalizations of the energy conditions to the averaged forms. As we have seen in Sec. 2.4, 2.5, and 2.6, the energy conditions play a key role mainly through the focusing theorems, we will focus on the averaged SEC and NEC. As a weaker condition than SEC, the averaged strong-energy condition (ASEC) is simply formulated as follows.
@@
  Definition 2.7.1 (Averaged SEC)   For every inextendible timelike geodesic curve with proper time $\tau$ and the tangent $\xi^{\mu}$, the integral along the timelike geodesic, $\int{\rm d}\tau\left(T_{\mu\nu}-\dfrac{1}{2}Tg_{\mu\nu}\right)\xi^{\mu}\xi^{\nu}\geqslant 0\,.$ (50)
@@
  Definition 2.7.2 (Averaged NEC)   For every inextendible null geodesic curve with affine parameter $\lambda$ and the tangent $k^{\mu}$, the integral along the null geodesic, $\int{\rm d}\lambda\,T_{\mu\nu}k^{\mu}k^{\nu}\geqslant 0\,.$ (52)